# Self-Rewarding Vision-Language Model via Reasoning Decomposition

### v1

@@ Abstract
Vision-Language Models (VLMs) often suffer from visual hallucinations – saying things that aren’t actually in the image – and language shortcuts, where they skip the visual part and just rely on text priors. These issues arise because most post-training methods for VLMs rely on simple verifiable answer matching and supervise only final outputs, leaving intermediate visual reasoning without explicit guidance. As a result, VLMs receive sparse visual signals and often learn to prioritize language-based reasoning over visual perception. To mitigate this, some existing methods add visual supervision using human annotations or distilled labels from external large models. However, human annotations are labor-intensive and costly, and because external signals cannot adapt to the evolving policy, they cause distributional shifts that can lead to reward hacking.
In this paper, we introduce Vision-SR1, a self-rewarding method that improves visual reasoning without relying on external visual supervisions via reinforcement learning. Vision-SR1 decomposes VLM reasoning into two stages: visual perception and language reasoning. The model is first prompted to produce self-contained visual perceptions that are sufficient to answer the question without referring back the input image. To validate this self-containment, the same VLM model is then re-prompted to perform language reasoning using only the generated perception as input to compute reward. This self-reward is combined with supervision on final outputs, providing a balanced training signal that strengthens both visual perception and language reasoning. Our experiments demonstrate that Vision-SR1 improves visual reasoning, mitigates visual hallucinations, and reduces reliance on language shortcuts across diverse vision-language tasks.
Figure: (Left) Case study comparing Vision-R1 with our method, showing reduced hallucinations. (Right) Reinforcement learning training curve compared to Vision-R1, along with accuracy on three widely used vision-language benchmarks.
@@ 1 Introduction
Recent advances in vision-language models (VLMs) have progressed by integrating pre-trained language models and vision encoders with instruction tuning (Liu et al., 2023b; et al, 2024; Chen et al., 2024; Bai et al., 2025; Li et al., 2025d). Despite these successes, a critical limitation remains in their reasoning capabilities: VLMs often produce visual hallucinations – descriptions of content that is not actually present in the image (Guan et al., 2024; Liu et al., 2024; Li et al., 2025e; Liu et al., 2023a) – or rely on language shortcuts, where the model bypasses visual understanding and instead depends solely on text priors (Si et al., 2022; Bleeker et al., 2024). Very recently, R1-style reinforcement learning (RL) methods have been shown to improve the reasoning abilities of VLMs across diverse tasks (Huang et al., 2025; Shen et al., 2025; Xia et al., 2025; Zhang et al., 2025). However, these methods often encourage “thinking over seeing” that lean heavily on language reasoning while underutilize visual perception (Liu et al., 2025; Yao et al., 2025). This imbalance makes VLMs susceptible to reward hacking (Fu et al., 2025) and spurious effects (Shao et al., 2025) observed in RL training. Although VLMs trained with RL often have apparent improvements, they can largely reflect probability shifts toward the style of training and test data, leading to language shortcut answers from prior knowledge and overlooking hallucination risks (Li et al., 2025b).
@@
In essence, most existing post-training methods for VLMs rely on simple verifiable answer matching and thus lack explicit supervision of visual information. As a result, VLMs receive sparse visual signals and often learn to prioritize language-based reasoning over visual perception. To mitigate this, some methods introduce intermediate visual supervision through human annotations (Thawakar et al., 2025) or distilled labels (e.g., pre-extracted key steps) from external models (Xu et al., 2024; Zhang et al., 2025; Xiao et al., 2025; Xia et al., 2025; Lu et al., 2025). However, these solutions face significant limitations. Human annotations are labor-intensive, costly, and difficult to scale across multimodal tasks, while distilled signals inherit biases from source models and often fail to generalize across diverse domains. Moreover, distributional shifts between fixed intermediate signals and the continually updated policy can lead to reward hacking (Gao et al., 2023). Most importantly, both approaches still rely on external supervision, restricting their scalability and applicability.
@@
In this paper, we introduce Vision-SR1, a reinforcement learning framework that encourages VLM to produce self-contained visual reasoning that can be verified by the VLM itself, without external supervision. Vision-SR1 decomposes the reasoning process into two stages: visual perception and language reasoning. The visual perception is required to capture all details relevant to answering the query, so that the reasoning stage can proceed without re-accessing the original image.
The training involves two rollout passes of the same VLM:
@@
  • An accuracy reward is computed by comparing the final answer with the ground truth.
@@
Both rewards are combined to provide a balanced training signal that strengthens both visual perception and language reasoning. Experiments show that Vision-SR1 improves visual reasoning, mitigates hallucinations, and reduces language shortcuts across diverse vision-language tasks.
@@ 2 Method
We build upon advancement of multimodal Group Relative Policy Optimization (GRPO) (Shao et al., 2024) for improving VLM reasoning. We first review the key concepts then introduce our method.
@@ 2.1 Preliminary: Reinforcement Learning for VLM with Verifiable Reward
We denote a pre-trained VLM as a policy model $\pi$ to be optimized in reinforcement learning. Given a multimodal question ($Q$) consists of an image $i$ and a text question $q$, where $Q=\{i,q\}$, the policy model $\pi$ generates a reasoning response $s$. We use GRPO to optimize the response $s$ for the policy model. For each multimodal question $Q=\{i,q\}$ we sample a group of $K$ candidate responses $\mathcal{S}_{Q}=\{s_{1},\dots,s_{K}\},\;s_{k}\sim\pi_{\theta}(\,\cdot\,|Q).$ Each response is scored by a scalar reward $r(Q,s_{k})$ (defined in Sec. 2.2), and we compute a group‑relative advantage
@@
which centres rewards within the group, removing question‑level biases while retaining pairwise preferences. We update the policy by maximizing
@@
The group‑centred baseline in (1) guarantees $\sum_{k}\hat{A}^{\text{grp}}(Q,s_{k})=0$, thereby reducing the variance of policy‑gradient estimates without requiring an external value critic.
2.2 Our Method: Self-Rewarding VLM via Reasoning Decomposition
As we discussed, incorporating intermediate visual supervision can strengthen the reasoning ability of VLMs. However, existing methods suffer from key limitations: methods based on human annotations are labor-intensive and costly (Thawakar et al., 2025), while approaches that distill supervision from external models provide static signals that cannot adapt as the policy model itself evolves during training (Zhang et al., 2025; Xiao et al., 2025; Xia et al., 2025). To overcome these issues, we introduce a self-rewarding framework that enables the VLM to reward its own visual perception. The key idea is to decompose the visual reasoning process into structured components, i.e., the VLM first produces a self-contained visual perception and then assesses whether this perception is sufficient for produce the final answer. This decomposition reduces reliance on external supervision and allows the reward signal to adapt dynamically as the model improves.
Decomposed VLM Reasoning.
To encourage the VLM to perform self-contained visual reasoning, we require every response to adhere to a See-Think generation format (Jia et al., 2024; Xia et al., 2025) format. Specifically, for a vision-language task, $Q=\{i,q\}$ where $i$ is the input image and $q$ is the textual query, the model produces the following structured output:
$\text{{<visual perception{}>}}\,c\;\text{{</visual perception{}>}}\;\|\;\text{{<think>}}\,t\;\text{{</think>}}\;\|\;\text{{<answer>}}\,a\;\text{{</answer>}},$
where $c$ is a self-contained visual perception that captures all visual information necessary to solve the task, so that the following language reasoning can proceed without re-accessing the original input image. Besides, $t$ is the language reasoning trace, and $a$ denotes the final answer.
Self-Reward for Visual Perception.
A challenge is judging whether the visual perception $c$ is self-contained – i.e. whether it encodes all the visual information needed to answer the question $Q=\{i,q\}$ correctly. To address this, our idea is to treat the visual perception as a text-only proxy for the image and validate it by re-prompting the VLM itself to perform language reasoning using only the generated perception as input. If the model can derive the correct answer from $(c,q)$ alone, we consider $c$ to be visually faithful and assign the corresponding visual reward.
@@ 2.2 Our Method: Self-Rewarding VLM via Reasoning Decomposition
where $a^{*}$ is the ground-truth answer. Instead of using an external reward model, we leverage the policy model’s own reasoning ability for self-evaluation. The model itself determines the reward by answering the question using only its generated visual perception (Figure 1).
Total Reward.
@@
The total reward comprises three aligned components, each conditioned on the question $Q=\{i,q\}$:
@@
  • Format reward $r_{\mathrm{fmt}}(s)$: measures whether the response strictly follows the required layout.
@@
  • Accuracy reward $r_{\mathrm{ans}}(Q,a)$: measures the correctness of the final answer. Because $a$ is generated after the reasoning trace $t$, the term implicitly rewards CoT reasoning.
@@
  • Visual reward $r_{\mathrm{cap}}(Q,c)$: measures whether the visual perception output is self-contained, i.e., sufficient to answer the question without image. A reward of 1 is assigned if, given only the question and the visual perception, the VLM can give the correct answer.
The entangling scalar reward is weighted as
$r(Q,s)=\,r_{\mathrm{visual}}(Q,c)+\,r_{\mathrm{ans}}(Q,a),+\alpha\,r_{\mathrm{fmt}}(s)$ (4)
with hyper-parameters $(0\leq\alpha\leq 1)$ tuned empirically. The answer accuracy reward $r_{\mathrm{ans}}(Q,a)$ is computed on the final answer $a$, produced autoregressively after the visual perception $c$. Because the decoder conditions on $c$, this term implicitly encourages coherent reasoning and tight coupling between visual perception and textual reasoning.
Figure 1: Overall framework of Vision-SR1. During RL training, the VLM performs two rollouts. In the first pass, the model takes an image–query pair and generates a structured output (visual perception, CoT reasoning, and answer), with answer reward computed against the ground truth. In the second pass, the model is re-prompted to answer using only query and its generated visual perception. If the correct answer is derived, a self-visual reward is assigned.

### v2

# Vision-SR1: Self-Rewarding Vision-Language Model via Reasoning Decomposition and Multi-Reward Policy Optimization

@@ Abstract
Vision-Language Models (VLMs) often suffer from visual hallucinations – generating things that are not consistent with visual inputs – and language shortcuts, where they skip the visual part and just rely on text priors. These issues arise because most post-training methods for VLMs rely on simple verifiable answer matching and supervise only final outputs, leaving intermediate visual reasoning without explicit guidance. As a result, VLMs receive sparse visual signals and often learn to prioritize language-based reasoning over visual perception. To mitigate this, some existing methods add visual supervision using human annotations or distilled labels from external large models. However, human annotations are labor-intensive and costly, and external signals can introduce high latency cost.
In this paper, we introduce Vision-SR1, a three-stage self-rewarding reinforcement learning method that improves visual reasoning without relying on external visual supervision. Vision-SR1 decomposes VLM reasoning into two components: visual reasoning and language reasoning, where the model is first prompted to produce self-contained visual descriptions sufficient to answer the question without referring back to the input image, before jointly optimizing both visual and language reasoning through our multi-reward loss objective. To validate this self-containment, the same VLM model is re-prompted to perform language reasoning using only the generated visual reasoning as input to compute visual reward. The final reward is computed through a decoupled reward-advantage framework, where visual reward and language reasoning reward each have their advantages, log probabilities, and KL divergence calculated separately. This decoupling enables more fine-grained reward computation by preventing the entanglement of heterogeneous reward signals. Our experiments show that Vision-SR1 improves visual reasoning, mitigates visual hallucinations, and reduces reliance on language shortcuts across diverse vision-language tasks, while being more efficient than methods that rely on external visual reward models, which require additional GPUs to host. In contrast, Vision-SR1 introduces no extra GPU overhead beyond that of standard training.
@@ 1 Introduction
Recent advances in vision-language models (VLMs) have progressed by integrating pre-trained language models and vision encoders with instruction tuning (Liu et al., 2023b; et al, 2024; Chen et al., 2024; Bai et al., 2025; Li et al., 2025d). Despite these successes, a critical limitation remains in their reasoning capabilities: VLMs often produce visual hallucinations—descriptions of content that is not actually present in the image (Guan et al., 2024; Liu et al., 2024; Li et al., 2025e; Liu et al., 2023a)—or rely on language shortcuts, where the model bypasses visual understanding and instead depends solely on text priors (Si et al., 2022; Bleeker et al., 2024). R1-style reinforcement learning (RL) methods have recently improved the reasoning abilities of VLMs across diverse tasks (Huang et al., 2025b; Shen et al., 2025; Xia et al., 2025; Zhang et al., 2025). However, these methods often encourage “thinking over seeing”, leaning heavily on language reasoning while demoting visual perception (Liu et al., 2025; Yao et al., 2025). This imbalance makes VLMs susceptible to reward hacking (Fu et al., 2025) and spurious effects (Shao et al., 2025) in RL training. Although VLMs trained with RL often “improve”, these improvements often just are probability shifts toward the style of training and test data, leading to language shortcut answers from prior knowledge and overlooking hallucination risks (Li et al., 2025b).
@@
In essence, most existing post-training methods for VLMs rely on a one stage, simple, verifiable answer matching and thus lack explicit supervision for visual information. As a result, VLM’s visual signals are sparse, leading them to prioritize language-based reasoning over visual perception. To mitigate this, some methods introduce intermediate visual supervision through human annotations (Thawakar et al., 2025) or distilled labels (e.g., pre-extracted key steps) from external models (Xu et al., 2024; Zhang et al., 2025; Xiao et al., 2025; Xia et al., 2025; Lu et al., 2025). However, these solutions have significant limitations. Human annotations are labor-intensive, costly, and difficult to scale across multimodal tasks, while distilled signals inherit biases and latency from source models and often fail to generalize across diverse domains. Moreover, distributional shifts between fixed intermediate signals and the continually updated policy can lead to reward hacking (Gao et al., 2023). Most importantly, both approaches remain limited by their reliance on external supervision or simply summimg up multiple intermediate rewards, restricting their scalability and applicability.
@@
This paper introduces Vision-SR1, a reinforcement learning framework that encourages VLMs to produce self-contained visual reasoning that can be verified by the VLM itself without external supervision. Vision-SR1 explicitly isolates the visual grounding stage from the reasoning stage: visual perception and language reasoning. The visual perception is required to capture all details relevant to answering the query, so that the reasoning stage can proceed without re-accessing the original image. We explicitly compute advantages and rollouts separately for each stage, then calculate individual Actor policy losses and KL divergence terms for the visual perception and language reasoning stages before combining them into a unified training objective.
The training has two rollout passes and one training objective optimization of the same VLM:
@@
  • An accuracy reward compares the final answer with the ground truth.
@@
– Multi-Reward Policy Optimization (objective optimization):
  • The multi-reward policy optimization enables the policy model to receive distinct feedback for visual reasoning quality and answer accuracy through separate advantage computations and rollout-specific loss terms.
The self-rewarding process eliminates the computational overhead of deploying additional reward models on separate GPUs, while the decoupled reward signals are combined through our multi-policy loss objective to provide balanced training that strengthens both visual perception and language reasoning without the entangled learning signals of traditional reward summation. Vision-SR1 improves visual reasoning, mitigates hallucinations, and reduces language shortcuts across diverse vision-language tasks.
@@ 2 Method
We build on Group Relative Policy Optimization (Shao et al., 2024, GRPO) for improving VLM reasoning. We first review the key concepts then introduce our method.
Figure 1: Overall framework of Vision-SR1. During RL training the VLM has two rollouts. In the first pass, the model takes an image–query pair and generates a structured output (visual perception, CoT reasoning, and answer), with answer reward computed against the ground truth. In the second pass, the model is re-prompted to answer using only query and its generated visual perception. If the correct answer is derived, a self-visual reward is assigned. We compute the advantages and log probabilities for each rollout for Multi-Reward Policy Optimization.
@@ 2.1 Preliminary: Reinforcement Learning for VLM with Verifiable Reward
We optimize a pre-trained VLM as a policy $\pi$ to be optimized in reinforcement learning. Given a multimodal question ($Q$) with an image $i$ and a text question $q$, where $Q=\{i,q\}$, the policy model $\pi$ generates a reasoning response $s$. GRPO optimizes the response $s$ for the policy model. For each multimodal question $Q=\{i,q\}$ we sample a group of $K$ candidate responses $\mathcal{S}_{Q}=\{s_{1},\dots,s_{K}\},\;s_{k}\sim\pi_{\theta}(\,\cdot\,|Q).$ Each response is scored by a scalar reward $r(Q,s_{k})$ (defined in Sec. 2.2), and we compute a group‑relative advantage
@@
which centers rewards within the group, removing question‑level biases while retaining pairwise preferences. We update the policy by maximizing
@@
The group‑centred baseline in equation 1 guarantees $\sum_{k}\hat{A}^{\text{grp}}(Q,s_{k})=0$, thereby reducing the variance of policy‑gradient estimates without requiring an external value critic.
2.2 Stage 1 & 2: Self-Rewarding VLM via Reasoning Decomposition
Incorporating intermediate visual supervision can strengthen the reasoning ability of VLMs. However, existing methods suffer from key limitations: approaches based on human annotations are labor-intensive and costly (Thawakar et al., 2025), while those that distill supervision from external models introduce additional compute overhead and training latency (Zhang et al., 2025; Xiao et al., 2025; Xia et al., 2025). To overcome these issues, we introduce a self-reward framework that enables the VLM to reward its own visual reasoning. The key idea is to decompose the visual reasoning process into structured components, i.e., the VLM first produces a self-contained visual perception and then assesses whether this perception is sufficient for produce the final answer. This decomposition reduces reliance on external supervision and allows the reward signal to adapt dynamically as the model improves.
Decomposing VLM Reasoning.
To encourage the VLM to reason about its visual input, we require every response to adhere to a See-Think generation format (Jia et al., 2024; Xia et al., 2025) format. Specifically, for a vision-language task, $Q=\{i,q\}$ where $i$ is the input image and $q$ is the textual query, the model produces the following structured output:
$\langle\text{visual\_reasoning}\rangle\,c\,\langle/\text{visual\_reasoning}\rangle\;\|\;\langle\text{think}\rangle\,t\,\langle/\text{think}\rangle\;\|\;\langle\text{answer}\rangle\,a\,\langle/\text{answer}\rangle$
where $c$ is a self-contained visual reasoning that captures all visual information necessary to solve the task, so that the following language reasoning can proceed without re-accessing the original input image. Besides, $t$ is the language reasoning trace, and $a$ is the final answer.
Self-Reward for Visual Reasoning.
A challenge is judging whether the visual reasoning $c$ is self-contained: i.e., whether it encodes all the visual information needed to answer the question $Q=\{i,q\}$ correctly. Our idea is to treat the visual perception as a text-only proxy for the image and validate it by re-prompting the VLM itself to perform language reasoning using only the generated perception as input. If the model can derive the correct answer from $(c,q)$ alone, we consider $c$ to be visually faithful and assign the corresponding visual reward.
@@ 2.2 Stage 1 & 2: Self-Rewarding VLM via Reasoning Decomposition
where $a^{*}$ is the ground-truth answer. Instead of using an external reward model, we leverage the policy model’s own reasoning ability for self-evaluation. The model itself determines the reward by answering the question using only its generated visual reasoning (Figure 1).
Reward Composition.
@@
The reward combines three aligned components, each conditioned on the question $Q=\{i,q\}$:
@@
$\bullet$ Format reward $r_{\mathrm{fmt}}(s)$: measures whether the response strictly follows the required layout. This reward is applied to both Visual reward and Accuracy reward.
@@
$\bullet$ Answer reward $r_{\mathrm{ans}}(Q,a)$: measures the correctness of the final answer ($r_{acc}$) plus the corresponding format reward. Because $a$ is generated after the reasoning trace $t$, the term implicitly rewards CoT reasoning. This is computed at first rollout with hyper-parameters $(0\leq\alpha\leq 1)$:
$r_{ans}(Q,a)=\,r_{\mathrm{acc}}(Q,a),+\alpha\,r_{\mathrm{fmt}}(s)$ (4)
@@
$\bullet$ Visual reward $r_{\mathrm{visual}}(Q,c)$: measures whether the visual reasoning output is self-contained, i.e., sufficient to answer the question without image ($r_{vis\_acc}$) plus corresponding format reward. A reward of 1 is assigned if, given only the question and the visual reasoning, the VLM can give the correct answer. This is computed at second rollout:
$r_{visual}(Q,a)=\,r_{\mathrm{vis\_acc}}(Q,a),+\alpha\,r_{\mathrm{fmt}}(s)$ (5)
2.3 Stage 3: Multi-Reward Optimization with Multi-Advantage Loss Computation.
Simply summing the visual reasoning reward and the final-answer accuracy reward could produce a sparse and entangled learning signal: the policy has little to tell which rollout was responsible. To disentangle visual reasoning and answer accuracy assignment, we keep the two rollouts—answer generation and visual reasoning—separate throughout the update. Each rollout receives its own log-probabilities, advantage, and KL term, and the gradients are combined only at the end. This turns the single multi-reward problem into two single-reward sub-problems that share parameters with individually optimized feedback.
Reward-Specific Log-Probability Tracking.
During sampling we cache the behavioral log probabilities for every token in each rollout:
$\log\pi_{\text{old}}^{(i)}\!\bigl(a_{\text{ans},t}\bigr),\;\log\pi_{\text{old}}^{(i)}\!\bigl(a_{\text{visual},t}\bigr),\;$
where $a_{\text{ans},t}$ is the action at step $t$ of the first rollout, and $a_{\text{visual},t}$ is the action (token) at step $t$ of the second rollout. At update time we compute the corresponding $\log\pi_{\theta}$ under the current parameters to compute the policy and KL losses.
Group-wise $z$-Score Advantage.
For each reward we follow GRPO to compute the advantage:
$A_{\text{ans}}^{(i)}\;=\;\frac{r_{\text{ans}}^{(i)}-\mu_{\text{ans}}}{\sigma_{\text{ans}}+\varepsilon},\qquad A_{\text{visual}}^{(i)}\;=\;\frac{r_{\text{visual}}^{(i)}-\mu_{\text{visual}}}{\sigma_{\text{visual}}+\varepsilon},$ (6)
with means and standard deviations $\mu_{\text{ans}}=\frac{1}{B}\sum_{i}r_{\text{ans}}^{(i)},\;\sigma_{\text{ans}}^{2}=\frac{1}{B}\sum_{i}\bigl(r_{\text{ans}}^{(i)}-\mu_{\text{ans}}\bigr)^{2}$, where $B$ is the rollout batch size (and analogously for the visual group). Broadcasting $A_{\text{ans}}$ to all caption tokens and $A_{\text{visual}}$ to all answer tokens gives two advantage masks that weight the corresponding log-probabilities during backpropagation for each sub-task.
Actor Loss (Policy Gradient Loss).
The actor loss computes weighted policy gradients for the two reward signals (answer and visual), with separate coefficients $\lambda_{\text{ans}}$ and $\lambda_{\text{visual}}$ indicating their contributions. [^1] We use 0.5 for $\lambda_{\text{ans}}$ and $\lambda_{\text{visual}}$.
$\mathcal{L}_{\text{actor}}=-\frac{1}{2B}\sum_{i,t}\left(A^{(i)}_{\text{ans},t}\,\log\pi_{\theta}(a^{(i)}_{\text{ans},t})+A^{(i)}_{\text{visual},t}\,\log\pi_{\theta}(a^{(i)}_{\text{visual},t})\right)$ (7)
, where $B$ is the rollout batch size (as in Eq. 6).
KL Divergence Regularization Loss.
The KL regularization applies separate penalty coefficients $\beta_{\text{cap}}$ and $\beta_{\text{ans}}$ to prevent excessive policy deviation for each reward component. $\mathcal{L}_{\text{KL}}$ $=\frac{\beta_{\text{ans}}}{B}\sum_{i=1}^{B}\sum_{t}\bigl[\log\pi_{\text{old}}\!\bigl(a^{(i)}_{\text{ans},t}\bigr)-\log\pi_{\theta}\!\bigl(a^{(i)}_{\text{ans},t}\bigr)\bigr]\;+\;\frac{\beta_{\text{visual}}}{B}\sum_{i=1}^{B}\sum_{t}\bigl[\log\pi_{\text{old}}\!\bigl(a^{(i)}_{\text{visual},t}\bigr)-\log\pi_{\theta}\!\bigl(a^{(i)}_{\text{visual},t}\bigr)\bigr$] (8)
Multi-Reward Loss Objective.
The total loss combines multi-reward policy gradients with component-specific regularization to optimize the model across both reward signals.
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{actor}}+\mathcal{L}_{\text{KL}}$ (9)